# Gravitational Wave Memory of Primordial Black Hole Mergers

### v1

@@ I Introduction
A major challenge in this field is to unravel the origin of binaries of compact objects. Such systems may arise either from astrophysical processes in the late universe or from primordial density fluctuations in the early universe, potentially leading to the formation of Primordial Black Holes (PBHs) Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx ; Carr:1975qj (see e.g. Byrnes:2025tji for recent reviews). PBHs, remnants of the early universe, offer an exciting opportunity to probe fundamental physics. They may also constitute a significant fraction of the universe’s Dark Matter (DM) content (see Carr:2020gox for a review of current constraints) and lead to detectable GW merger events Bird:2016dcv ; Clesse:2016vqa ; Sasaki:2016jop ; Eroshenko:2016hmn ; Wang:2016ana ; Ali-Haimoud:2017rtz ; Chen:2018czv ; Raidal:2018bbj ; Franciolini:2021tla ; Franciolini:2021nvv .
@@
In this situation, we focus on the low-frequency tail of the GW signal of light PBH binaries, which may be probed by the LVK network or upcoming GW detectors such as the Einstein Telescope (ET) Franciolini:2023opt ; Abac:2025saz and LISA LISACosmologyWorkingGroup:2023njw ; LISA:2024hlh . PBH binaries generate a GW signal at frequencies far below the peak merger signal at two distinct phases of their evolution. First, during their early inspiral, they emit quasi-monochromatic GWs. Second, the merger produces a characteristic low-frequency GW signature known as the GW memory effect Christodoulou:1991cr ; Blanchet:1992br ; Thorne:1992sdb (see Mitman:2024uss for a recent review).
@@
Searches for both the primary and memory signal benefit significantly from the implementation of a matched filtering search, i.e. the use of a signal template. For the primary signal this is a well-established technique maggiore2008gravitational . For the memory signal, the properties of the signal templates depend crucially on the detector response function, see Ref. Ghosh:2023rbe for LISA and Ref. Ebersold:2020zah for an approximation of memory templates for LIGO. Remarkably, the resulting waveform can be to good approximation characterized by very few parameters. We revisit and refine these derivations, including the full instrument response function and discussing the implications for a matched filtering search.
The structure of the paper is as follows: In Section II, we review the GW memory effect associated with BH mergers, modeling the GW signal and comparing it with the inspiral GW emission. Section III discusses the PBH merger rates used in our analysis. Section IV details our methodology and results before we conclude in Section V. The appendices contain: A) a derivation of the GW memory signal; B) a derivation of the memory signal templates at interferometers; C) the sensitivity curves of the detectors considered in this paper. In what follows we set $c=G=1$.
@@ II.1 GW memory signal and modeling
The memory effect is intimately connected to BMS symmetries, which describe the symmetries of asymptotic space-times in General Relativity (GR). Recently, this has also been referred to as displacement memory because related to the supertranslation symmetry and it represents the dominant component of new memories linked to different symmetries which leave the asymptotic space-time invariant Flanagan:2018yzh ; Grant:2021hga ; Grant:2022bla ; Siddhant:2024nft ; Pasterski:2015tva ; Nichols:2018qac ; DeLuca:2024bpt . Furthermore, it is tied to the quantum structure of scattering amplitudes through soft theorems Strominger:2014pwa ; Strominger:2017zoo . The detection of the GW memory effect is not only an additional confirmation of GR, but it is also relevant for its connection with the infrared structure of gravity and gauge theories. Indeed, asymptotic symmetries, soft theorems and memories represent the three different corners of the triangle diagram describing the infrared structure of gauge theories Strominger:2014pwa ; Strominger:2017zoo .
When solving for the gravitational wave (GW) sourced by its energy-momentum tensor, the memory effect can be computed as an integral over the energy flux of the primary GWs Wiseman:1991ss ; Thorne:1992sdb ; Favata:2010zu ; Heisenberg:2023prj :
@@
$\frac{\mathrm{d}E}{\mathrm{d}u^{\prime}\mathrm{d}\Omega^{\prime}}=\frac{d^{2}c^{3}}{16\pi}\langle\dot{h}_{+}^{2}+\dot{h}_{\times}^{2}\rangle,$ (2)
where $h_{+}$ and $h_{\times}$ are the amplitudes of the corresponding polarization states and the spacetime average delimited-⟨$\langle$..$\rangle$ … ⟩ is to ensure a well-defined energy content of the GWs PhysRev.121.1556 ; PhysRev.166.1272 ; maggiore2008gravitational ; Favata:2011qi ; Heisenberg:2023prj .
Decomposing the GW strain, $h=h_{+}-ih_{\times}$ in spin-weighted spherical harmonics, as defined in Eq. (30) in terms of $(\ell,m)$, the energy flux $\propto|\dot{h}_{(2,2)}|^{2}$ is largely dominated by the $(2,2)$ mode. Due to symmetry, this means the memory primarily appears in the $(2,0)$ mode, which in the reference frame of the source with the $z$-axis pointing in the direction of the observer, can be identified as the $+$ polarization, as explained in App. A.
Figure 1: Comparison between the dominant oscillatory gravitational waveform, $h_{\rm main}$, associated with the (2,2) mode, and the memory waveform, $h_{\rm mem}$, for an edge-on system. For reference, we also include the analytical approximation of the memory from Eq. (3), using a characteristic rise time of $\Delta\tau=13M$.
A key feature of GW memory is that its value at any given time depends on the entire history of the binary—making it a hereditary effect. This results in a final amplitude comparable to the peak of the dominant oscillatory signal. From numerical waveforms,  [^1] We compute the memory from the pubic package GWMemory PhysRevD.98.064031 which numerically evaluates Eq. (1) for given input waveforms. For the latter, we used the NRHybSur3dq8 waveform model PhysRevD.99.064045 . we find that the peak strain satisfies $\max\{h\}\simeq 3.7\max\{h_{\rm mem}\}$ for an edge-on system, where the memory is maximized. Fig. 1 shows both the dominant oscillatory signal, $h_{\rm main}$, which is the $(2,2)$ mode and the memory $h_{\rm mem}$, which is the $(2,0)$ mode. The latter follows a characteristic step-like behavior, with the transition occurring around merger time $t_{\rm m}$, when most of the energy is released.
@@
$h_{\rm tanh}(t)=\frac{\Delta h_{\rm mem}}{2}\left[\tanh{\left(\frac{-\tau}{\Delta\tau}\right)}+1\right],$ (3)
where $t_{m}-t\equiv\tau$ is the time to merger and where $\Delta\tau$ defines the rise time of the memory and $\Delta h_{\rm mem}$ its final amplitude. We find that the difference between the final value and the initial value of the ${(2,0)}$ mode is $d/M\Delta h_{2,0}\simeq 0.08$ for a system of equal mass binary with total mass $M$ at distance $d$, which leads to 222 This is somewhat lower than the value given in Ref. Pollney_2011 , which obtained $d/Mh_{2,0}=0.097\pm 0.002$. This can be traced back to the initial time inserted as a lower integration boundary in Eq. (1), which in Ref. Pollney_2011 was taken to be $\infty$- ∞ whereas in this paper we focus on the memory produced around the merger time. Concretely, we consider an initial time $\tau_{\rm in}=10^{5}M$, capturing roughly the final 100 cycles, as shown in Fig. 1. Our results are robust to varying this choice.
@@ II.2 Memory signal template
As discussed in the previous section, due to the characteristic low-frequency behavior of the memory, it is, in principle, possible to detect a merger occurring at frequencies above the detector’s sensitivity range. To our knowledge, the only searches for out-of-band mergers using real data have been conducted in Ref. Agazie:2025oug (see also NANOGrav:2019vto ; NANOGrav:2023vfo ) in the context of pulsar timing array GW searches, and the authors of Ebersold:2020zah , who searched for subsolar mergers in the second observing run (O2) of the LIGO and Virgo detectors. Using LIGO/Virgo as triggers to search of out-of-band memory signals in LISA was proposed in Ghosh:2023rbe . In this regime, the rise time of the memory corresponds to frequencies higher than the sensitivity range of the instrument, which is primarily sensitive to the $f^{-1}$ component of the signal. Modeling the detector as a high-pass filter motivated by the rapid drop-off in sensitivity below a frequency $f_{\rm min}^{\rm det}$, Ref. Ebersold:2020zah employed the unmodeled search method coherent WaveBursts (cWB), which, while efficient and model-independent, is not as sensitive as a matched filtering search.  [^4] By comparing the sensitivity of the cWB search with our estimates, using updated sensitivity curves, we found that the latter are about two orders of magnitudes less sensitive. Given the simplicity and universality of the memory waveform in this regime, we argue that incorporating a memory-based template search into current and future analyses is not only feasible but essential because it allows probing parts of the parameter space that otherwise remain unprobed.
@@
[…]
Here ${\bm{k}}=2\pi f\hat{\bm{k}}$ denotes the wave vector of the GW in the detector frame for a GW originating from a direction $-{\hat{\bm{k}}}$ in the sky. Assuming that the primary signal peaks significantly out-of band, $f\ll\Delta\tau^{-1}$, and that we can model the detector response as a flat response in frequency limited to a frequency window $[f_{\rm min}^{\rm det},f_{\rm max}^{\rm det}]$, inserting (II.1) yields,
[…]
@@
To illustrate this, Fig. 3 compares the leading order terms of Eq. (8) [black] with the full result in Eq. (6) [colored] for the case of an equal-mass binary with total mass $M=10^{-2}M_{\odot}$ at a distance of $d=1$ kpc in LIGO, assuming $f_{\rm min}^{\rm det}=10$ Hz, as in Ref. Ebersold:2020zah . Accounting for the full response function reveals that the time scale of the characteristic rapid change in the signal at $t=t_{m}$ is set by the effective interferometer arm length, which is about 4 ms for LIGO. (For the case of $f\ll\Delta\tau^{-1}$, otherwise the latter time scale would also enter.) This can be understood intuitively as the memory signal Eq. (3) changing the effective interferometer armlength. The slower oscillations are induced by the high-pass filter in Eq. (II.2), and consequently, their frequency is set by $f_{\rm min}^{\rm det}$. The latter is given by either the (uninterrupted) observation time of the detector or the high-pass filters applied in the analysis (as done in particular in most LIGO analyses to remove the frequency range dominated by seismic noise). 555 We caution that the oscillatory features associated with $f_{\rm min}^{\rm det}$ are induced by the data processing and hence by construction are present in the data even in the absence of a memory signal. They should hence not be used for signal search. One such period, corresponding to $0.11/f_{\rm min}^{\rm det}=$ 0.1 s, is highlighted by the dashed gray vertical lines in Fig. 3. The amplitude of the band passed memory is exponentially suppressed as $e^{-2\pi f_{\rm min}^{\rm det}t}$ as shown with the purple dotted lines.
@@ II.3 GWs signals from inspiraling binaries
with $Q_{+}(\iota)=(1+\cos^{2}\iota)/2$ and $Q_{\times}=\cos\iota$; $G$ is Newton’s constant; ${\cal M}\equiv(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass of a binary with constituent masses $m_{1}$, $m_{2}$; $d$ is the proper distance from the binary to the observer and $\psi$ denotes the phase of the binary. We assume the GW emitted by the inspiral is described by Eq. (9) up to the Innermost Stable Circular Orbit (ISCO) Maggiore:2007ulw
$f_{\rm ISCO}=2.2{\rm kHz}\left(\frac{M_{\odot}}{m_{1}+m_{2}}\right).$ (10)
@@ III.1 PBH merger rates
×$\times\left[\frac{m_{1}m_{2}}{(m_{1}+m_{2})^{2}}\right]^{-\frac{34}{37}}\left(\frac{m_{1}+m_{2}}{10^{-12}M_{\odot}}\right)^{-\frac{32}{37}}$
×$\times S\left(t,M,f_{\rm PBH},\psi\right)\psi(m_{1})\psi(m_{2}),$ (17)
@@
The rate shown in Eq. (17) is boosted when considering small distances from the Earth due to the galactic overdensity. Assuming the distribution of PBH binaries to follow the one of the dark matter, the local merger rate density is enhanced by the overall factor $[1+\delta(d)]$ included in Eq. (17) Pujolas:2021yaw . This gives the differential merger rate density dependence on the distance from the earth $d$. We defined the overdensity factor $\delta(r)\equiv\rho_{\text{\tiny DM}}(r)/\bar{\rho}_{\text{\tiny DM}}$. The Milky Way dark matter halo is modeled as a Navarro-Frenk-White density profile, $\rho_{\text{\tiny{DM}}}(r)={\rho_{0}}r_{0}/[r\left(1+{r}/{r_{0}}\right)^{\!2}]\,$, with the normalization $\rho_{0}$ fixed such that $\rho_{\text{\tiny DM}}(r=r_{\odot})=7.9\times 10^{-3}M_{\odot}/{\rm pc}^{3}$ at the the solar system location is $r_{\odot}\simeq 8.0\,{\rm kpc}$, and $r_{0}=15.6\,{\rm kpc}$. We can simplify the overdensity within a volume of radius $d$ around the Earth location as
@@
In conclusion, while these alternative binary formation channels are possible, they remain subdominant within the standard scenario Ali-Haimoud:2017rtz ; Franciolini:2022ewd ; Raidal:2024bmm . This is particularly true for light PBHs with masses below the stellar range, as the respective event rates scale as follows: ${\cal R}^{\text{EU}}\propto m^{-32/37},{\cal R}^{\text{cap}}\propto m^{-11/21},{\cal R}^{\text{3b}}\propto m^{-(11\div 16)/21}.$

### v2

@@ I Introduction
A major challenge in this field is to unravel the origin of binaries of compact objects. Such systems may arise either from astrophysical processes in the late universe or from primordial density fluctuations in the early universe, potentially leading to the formation of Primordial Black Holes (PBHs) Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx ; Carr:1975qj (see e.g. Byrnes:2025tji for recent reviews). PBHs, remnants of the early universe, offer an exciting opportunity to probe fundamental physics. They may also constitute a significant fraction of the universe’s Dark Matter (DM) content (see Carr:2020gox for a review of current constraints) and lead to detectable GW merger events Bird:2016dcv ; Clesse:2016vqa ; Sasaki:2016jop ; Eroshenko:2016hmn ; Wang:2016ana ; Ali-Haimoud:2017rtz ; Chen:2018czv ; Raidal:2018bbj ; Franciolini:2021tla ; Franciolini:2021nvv ; Afroz:2024fzp .
@@
In this situation, we focus on the low-frequency tail of the GW signal of light PBH binaries, which may be probed by the LVK network or upcoming GW detectors such as the Einstein Telescope (ET) Franciolini:2023opt ; Abac:2025saz and LISA LISACosmologyWorkingGroup:2023njw ; LISA:2024hlh . PBH binaries generate a GW signal at frequencies far below the peak merger signal at two distinct phases of their evolution. First, during their early inspiral, they emit quasi-monochromatic GWs. Second, another low-frequency GW signature, known as the GW memory Christodoulou:1991cr ; Blanchet:1992br ; Thorne:1992sdb (see Mitman:2024uss for a recent review), is generated by all GW sources and reaches its maximum strength during the merger phase for a binary coalescence.
@@
Searches for both the primary and memory signal benefit from the implementation of a matched filtering search, i.e. the use of a signal template. For the primary signal this is a well-established technique maggiore2008gravitational . For the memory signal, the properties of the signal templates depend crucially on the detector response function, see Ref. Ghosh:2023rbe for LISA and Ref. Ebersold:2020zah for an approximation of memory templates for LIGO. Remarkably, the resulting waveform can be to good approximation characterized by very few parameters. We revisit and refine these derivations, including the full instrument response function and discussing the implications for a matched filtering search.
The structure of the paper is as follows: In Section˜II, we review the GW memory effect associated with BH mergers, modeling the GW signal and comparing it with the inspiral GW emission. Section III discusses the PBH merger rates used in our analysis. Section IV details our methodology and results before we conclude in Section˜V. The appendices contain: A) a derivation of the GW memory signal; B) a derivation of the memory signal templates at interferometers; C) the sensitivity curves of the detectors considered in this paper. In what follows we set $c=G=1$.
@@ II.1 GW memory signal and modeling
The memory effect is intimately connected to BMS symmetries, which describe the symmetries of asymptotic space-times in General Relativity (GR). In particular, the conservation law associated with supertranslation symmetry provides an alternative formulation for computing the GW memory compared to the one presented here and in more detail in App. A. Because this effect leaves a permanent displacement between two inertial observers, it is commonly referred to in recent literature as the displacement memory. However, this effect has been shown to represent only the leading contribution among several distinct types of memory, each linked to different asymptotic symmetries that preserve the structure of spacetime at null infinity Flanagan:2018yzh ; Grant:2021hga ; Grant:2022bla ; Siddhant:2024nft ; Pasterski:2015tva ; Nichols:2018qac ; DeLuca:2024bpt . Furthermore, it is tied to the quantum structure of scattering amplitudes through soft theorems Strominger:2014pwa ; Strominger:2017zoo . The detection of the GW memory effect is not only an additional confirmation of GR, but it is also relevant for its connection with the infrared structure of gravity and gauge theories. Indeed, asymptotic symmetries, soft theorems and memories represent the three different corners of the triangle diagram describing the infrared structure of gauge theories Strominger:2014pwa ; Strominger:2017zoo .
When solving for the GW sourced by its energy-momentum tensor, the memory effect can be computed as an integral over the energy flux of the primary GWs Wiseman:1991ss ; Thorne:1992sdb ; Favata:2010zu ; Heisenberg:2023prj :
@@
$\frac{\mathrm{d}E}{\mathrm{d}u^{\prime}\mathrm{d}\Omega^{\prime}}=\frac{d^{2}}{16\pi}\langle\dot{h}_{+}^{2}+\dot{h}_{\times}^{2}\rangle,$ (2)
where $h_{+}$ and $h_{\times}$ are the amplitudes of the corresponding polarization states and the spacetime average $\langle$$\rangle$ is to ensure a well-defined energy content of the GWs PhysRev.121.1556 ; PhysRev.166.1272 ; maggiore2008gravitational ; Favata:2011qi ; Heisenberg:2023prj .
Decomposing the GW strain, $h=h_{+}-ih_{\times}$ in spin-weighted spherical harmonics, as defined in Eq. (30) in terms of $(\ell,m)$, the energy flux $\propto|\dot{h}_{(2,2)}|^{2}$ is largely dominated by the $(2,2)$ mode. Due to symmetry, this means the memory primarily appears in the $(2,0)$ mode, which in the reference frame of the source with the $z$-axis pointing in the direction of the observer, can be identified as the ++ polarization, as explained in App. A.
Figure 1: Comparison between the dominant oscillatory gravitational waveform $({\rm main})$ associated with the (2,2) mode, and the memory waveform $({\rm mem})$ for an edge-on system. For reference, we also include the analytical approximation of the memory from Eq. (3), using a characteristic rise time of $\Delta\tau=13M$.
A key feature of GW memory is that its value at any given time depends on the entire history of the binary—making it a hereditary effect. This results in a final amplitude comparable to the peak of the dominant oscillatory signal. From numerical waveforms, [^1] We compute the memory from the public package GWMemory PhysRevD.98.064031 which numerically evaluates Eq. (1) for given input waveforms. For the latter, we used the NRHybSur3dq8 waveform model PhysRevD.99.064045 . we find that the peak strain satisfies $\max\{h\}\simeq 3.7\max\{h_{\rm mem}\}$ for an edge-on system, where the memory is maximized. Fig. 1 shows both the dominant oscillatory signal, $h_{\rm main}$, which is the $(2,2)$ mode and the memory $h_{\rm mem}$, which is the $(2,0)$ mode. The latter follows a characteristic step-like behavior, with the transition occurring around merger time $t_{\rm m}$, when most of the energy is released.
@@
[…]
where $t_{m}-t\equiv\tau$ is the time to merger and where $\Delta\tau$ defines the rise time of the memory and $\Delta h_{\rm mem}$ its final amplitude. We find that the difference between the final value and the initial value of the (2,0){(2,0)} mode is $d/M\Delta h_{2,0}\simeq 0.08$ for a system of equal mass binary with total mass $M$ at distance $d$, which leads to [^2] This is somewhat lower than the value given in Ref. Pollney_2011 , which obtained $d/Mh_{2,0}=0.097\pm 0.002$. This can be traced back to the initial time inserted as a lower integration boundary in Eq. (1), which in Ref. Pollney_2011 was taken to be -$\infty$ whereas in this paper we focus on the memory produced around the merger time. Concretely, we consider an initial time $\tau_{\rm in}=10^{5}M$, capturing roughly the final 100 cycles, as shown in Fig. 1. Our results are robust to varying this choice.
@@ II.2 Memory signal template
As discussed in the previous section, due to the characteristic low-frequency behavior of the memory, it is, in principle, possible to detect a merger occurring at frequencies above the detector’s sensitivity range. To our knowledge, the only searches for out-of-band mergers using real data have been conducted in Ref. Agazie:2025oug (see also NANOGrav:2019vto ; NANOGrav:2023vfo ) in the context of pulsar timing array GW searches, and the authors of Ebersold:2020zah , who searched for subsolar mergers in the second observing run (O2) of the LIGO and Virgo detectors. Using LIGO/Virgo as triggers to search for out-of-band memory signals in LISA was proposed in Ghosh:2023rbe . In this regime, the rise time of the memory corresponds to frequencies higher than the sensitivity range of the instrument, which is primarily sensitive to the $f^{-1}$ component of the signal. Modeling the detector as a high-pass filter motivated by the rapid drop-off in sensitivity below a frequency $f_{\rm min}^{\rm det}$, Ref. Ebersold:2020zah employed the unmodeled search method coherent WaveBursts (cWB) Klimenko:2008fu ; Klimenko:2015ypf , which, while efficient and model-independent, is in general not as sensitive as an optimal matched filtering search. [^4] By comparing the sensitivity of the cWB search in Ref. Ebersold:2020zah with our estimates, we find a potential sensitivity gain of two orders of magnitude. There are however caveats to this comparison, most notably that an actual search on real data faces challenges which are not accounted for in theoretical sensitivity estimates. Moreover, Ref. Ebersold:2020zah was performed on O2 data whereas our estimate is based on design sensitivity, see App. C. Given the simplicity and universality of the memory waveform in this regime, we argue that incorporating a memory-based template search into current and future analyses is not only feasible but essential because it allows probing weak events which otherwise would not pass the detection threshold of an unmodeled search.
@@
$s(t)=\int_{-\infty}^{\infty}dfR(\bm{k})\tilde{h}(f)e^{2\pi ift}\,.$ (6)
Here $^{\bm{k}}=2\pi f\hat{\bm{k}}$ denotes the wave vector of the GW in the detector frame for a GW originating from a direction $\hat{\bm{k}}$} in the sky. Assuming that the primary signal peaks significantly out-of band, $f\ll\Delta\tau^{-1}$, and that we can model the detector response as a flat response in frequency limited to a frequency window $f_{\rm min}^{\rm det},f_{\rm max}^{\rm det}]$, inserting (II.1) yields,
$s(t)$ $\simeq R(\hat{\bm{k}})\Delta h_{\text{mem}}\left(\int_{f_{\rm min}^{\rm det}}^{f_{\rm max}^{\rm det}}\frac{(e^{2\pi ift}-e^{-2\pi ift})}{2\pi if}df+\frac{1}{2}\right)$
@@
To illustrate this, Fig. 3 compares the leading order terms of Eq. (8) [black] with the full result in Eq. (6) [colored] for the case of an equal-mass binary with total mass $M=10^{-2}M_{\odot}$ at a distance of $d=1$ kpc in LIGO, assuming $f_{\rm min}^{\rm det}=10$ Hz, as in Ref. Ebersold:2020zah . Accounting for the full response function reveals that the time scale of the characteristic rapid change in the signal at $t=t_{m}$ is set by the effective interferometer arm length, which is about 4 ms for LIGO. (For the case of $f\ll\Delta\tau^{-1}$, otherwise the latter time scale would also enter.) This can be understood intuitively as the memory signal Eq. (3) changing the effective interferometer arm length. The slower oscillations are induced by the high-pass filter in Eq. (II.2), and consequently, their frequency is set by $f_{\rm min}^{\rm det}$. The latter is given by either the (uninterrupted) observation time of the detector or the high-pass filters applied in the analysis (as done in particular in most LIGO analyses to remove the frequency range dominated by seismic noise). [^5] We caution that the oscillatory features associated with $f_{\rm min}^{\rm det}$ are induced by the data processing and hence by construction are present in the data even in the absence of a memory signal. They should hence not be used for signal search. One such period, corresponding to $1/f_{\rm min}^{\rm det}=0.1$ s, is highlighted by the dashed gray vertical lines in Fig. 3. The amplitude of the band passed memory is exponentially suppressed as $e^{-2\pi f_{\rm min}^{\rm det}t}$ as shown with the purple dotted lines.
@@ II.3 GWs signals from inspiraling binaries
with $Q_{+}(\iota)=(1+\cos^{2}\iota)/2$ and $Q_{\times}=\cos\iota$; $G$ is Newton’s constant; ${\cal M}\equiv(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass of a binary with constituent masses $m_{1}$, $m_{2}$; $d$ is the proper distance from the binary to the observer and $\psi$ denotes the phase of the binary. We assume the GW emitted by the inspiral (being twice the orbital frequency) is described by Eq. (9) up to the Innermost Stable Circular Orbit (ISCO) Maggiore:2007ulw
$f_{\rm ISCO}=4.4{\rm kHz}\left(\frac{M_{\odot}}{m_{1}+m_{2}}\right).$ (10)
@@ III.1 PBH merger rates
$\times\left[\frac{m_{1}m_{2}}{(m_{1}+m_{2})^{2}}\right]^{-\frac{34}{37}}\left(\frac{m_{1}+m_{2}}{10^{-12}M_{\odot}}\right)^{-\frac{32}{37}}$
$\times S\left(t,M,f_{\rm PBH},\psi\right)\psi(m_{1})\psi(m_{2}),$ (17)
@@
The rate shown in Eq. (17) is boosted when considering small distances from the Earth due to the galactic overdensity. Assuming the distribution of PBH binaries to follow the one of the dark matter, the local merger rate density is enhanced by the overall factor $1+\delta(d)]$ included in Eq. (17) Pujolas:2021yaw . This gives the differential merger rate density dependence on the distance from the earth $d$. We defined the overdensity factor $\delta(r)\equiv\rho_{\text{\tiny DM}}(r)/\bar{\rho}_{\text{\tiny DM}}$. The Milky Way dark matter halo is modeled as a Navarro-Frenk-White density profile, $\rho_{\text{\tiny{DM}}}(r)={\rho_{0}}r_{0}/[r\left(1+{r}/{r_{0}}\right)^{\!2}]\,$, with the normalization $\rho_{0}$ fixed such that $\rho_{\text{\tiny DM}}(r=r_{\odot})=7.9\times 10^{-3}M_{\odot}/{\rm pc}^{3}$ at the the solar system location is $r_{\odot}\simeq 8.0\,{\rm kpc}$, and $r_{0}=15.6\,{\rm kpc}$. We can simplify the overdensity within a volume of radius $d$ around the Earth location as
@@
In conclusion, while these alternative binary formation channels are possible, they remain subdominant within the standard scenario Ali-Haimoud:2017rtz ; Franciolini:2022ewd ; Raidal:2024bmm . This is particularly true for light PBHs with masses below the stellar range, as the respective event rates scale as follows: ${\cal R}^{\text{EU}}\propto m^{-32/37},{\cal R}^{\text{cap}}\propto m^{-11/21},{\cal R}^{\text{3b}}\propto m^{-(11--16)/21}.$